OpenStudy (anonymous):

PLEASE HELP! PHYSICS TEST TOMORROW AND SO CONFUSED! ANSWERS INCLUDED PLEASE JUST SHOW HOW TO DO IT! A model rocket is lauched straight upward with an initial speed of 50.0 m/s. It accelerates with a constant upward acceleration of 2.00 m/s^2 until its engine stops at an altitude of 150 m. a. What is the maximum height reached by the rocket? (Answer= 310) b. When does the rocket reach maximum height? (Answer= 8.5) c. How long is the rocket in the air? (Answer= 16.4)

OpenStudy (anonymous):

ok we have 2 instances 1st instance, when the rocket is launched x_0=0 v_0=50 a=2 2nd instance, after the rocket reaches 150 m and its engine stops this means once it reaches 150m, acceleration is changes to -9.8m/s^2 which is the acceleration of gravity x_0=150 v_0=? --solved from 1st instance information a=-9.8

OpenStudy (anonymous):

okayy, and then what?

OpenStudy (anonymous):

well, you need to realize this is a 2 step process problem to find max height, you need to first determine the velocity the rocket reaches once it reaches 150m v_f=a(t)+v_o then v_f of first instance is v_o of 2nd instance using y=1/2 at^2+v_ot+y_o find the derivative and set y'=0 and solve for time then use that time and plug back in into y=1/2at^2... to find the max height it reaches

OpenStudy (anonymous):

for b. you first need to determine the time it took for the rocket to reach 150m in instance 1 then add that to the time it took to reach max height that was already determine in part a

OpenStudy (anonymous):

to determine how long the rocket is in the air, you need to determine the time when the rocket hits the ground using the 2nd instance information then add that to the time it took the rocket to reach 150m in the first instance to get total time in the air

OpenStudy (anonymous):

For A i got 55.68

OpenStudy (anonymous):

can i see your work?

OpenStudy (anonymous):

it would be impossible to get 55.68 for a because you know the rocket already reached 150m and was still travelling upwards due to the previous momentum/ veloccity of the rocket

OpenStudy (anonymous):

unless you were refering to the initial velocity for the 2nd instance

OpenStudy (anonymous):

idk what i did, hold on, and after can u show me the work cause im still pretty confused